- Data distribution
- Using the Archive
- Gemini Home
- Telescopes and Sites
- Science Visitors at Gemini
- Observing With Gemini
- Retired Instruments
- Visiting Instrument Policy
- Visiting Instrument Telescope Interfaces
- DSSI Speckle Camera
- TEXES (North)
- Integration Time Calculators
- Magnitudes and Fluxes
- Near-IR Resources
- Mid-IR Resources
- Observing Condition Constraints
- Performance Monitoring
- SV/Demo Science
- Future Instrumentation & Current Development
- Queue and Schedules
- Data and Results
- Calibration Data Retrieval
- Past/Future Development
Change page style:
Progress with NIRI during semester 2001B
30 Jan 2002
The 2001B semester has come to a close. Although no science observations could be completed, we have made progress towards our goal of starting regular science operations with NIRI.
Earlier in the semester UH installed new bearings on the beam steering mirrors that were responsible for the large image motions that hampered observations in 2001A. These new bearing were tested in NIRI on the telescope in September, and flexure within the science channel shown to be acceptable. We then dealt with additional mechanical problems and fixed a couple of seized bushings. By the end of November we were back on the sky and performing final acceptance tests. Weather conditions were not good enough to proceed with system verification or science observations, but we did complete some tests of new software versions and the NIRI mechanisms. Flexure between the science channel and the on-instrument wavefront sensor is still larger than the specification, but will not impact most NIRI programs. Further measurements during better weather conditions will help us understand how best to deal with the remaining flexure.
Science observations and additional commissioning work were scheduled for Dec. 28 through Jan. 7 and again Jan. 17 through 29, 2002. We were able to complete a number of basic engineering and software tests in spite of poor conditions. Among other things, we demonstrated the ability to fully sequence NIRI and the telescope to execute observations defined in the Observing Tool efficiently and automatically.
No science observations were completed during 2001B. The total time during the semester successfully used for science observations was 2 hours. The last two runs (23 nights) were scheduled for commissioning, system verification, and science queue programs. Weather conditions during the last month have been poor on Mauna Kea, and we lost almost all the time to wind, freezing fog, and snow. Only limited engineering work could be completed.
NIRI has been working well throughout the last two months, and is ready to go for science observations. We are optimistic that we will be able to perform system verification and queue observations as soon as the weather cooperates. We appreciate the efforts and patience of all the observers who dutifully prepared their 2001B Phase II proposals and worked hard to prepare their programs.
Thanks once again for your patience. Please submit any questions you may have through the Help Desk so that information can be appropriately distributed.
Last update Jan. 30, 2002; Joe Jensen